40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 8480     Entry time: Tue Apr 23 22:59:05 2013
Author: rana 
Type: Configuration 
Category: SUS 
Subject: Optical Lever Gains normalized 

Due to the recent addition of cal factors in the OL error points, the OLPIT_GAIN and OLYAW_GAIN have been reduce to tiny numbers (e.g. 0.002).

Since our MEDM only shows 3 digits past the decimal point by default, it makes more sense to have the gains around 1.

So I reduced the gains in all of the FM1 filters from 1000 to 1 and multiplied the GAIN values by 1000 (using ezcastep) to compensate.

All of the active optics seem to be behaving as before. Haven't tested ETMs or SRM yet.

ELOG V3.1.3-